LIST OF PUBLICATIONS BY Domokos SZÁSZ

1. On the general branching process with continuous time parameter. Studia Sci. Math. Hung. 2(1967), 227-246.

2. Spreading processes (in Hungarian). Graduate thesis. Budapest, 1967.

3. The applications of distribution functions in water resource management (in Hungarian). Hidrologiai közlöny, 1968, 433-446, (with M. Domokos)

4. Generation of fitting distribution functions of discharges by electronic computer. Publ. 81. IASH “The use of analogue and digital computers in hydrology” (Tuscon, Arizona, 1968), Vol. II, 535-545. (with M. Domokos)

5. Probability, Mathematical Statistics and their Applications. Lecture Notes. Ed.: Medgyessy and G. Tusnády, Math. Inst. of Hung. Acad. of Sciences, Budapest, 1968 (co-author)

6. Asymptotically uniform sequences of measures. Studia Sci. Math. Hung. 4(1969), 313-329.

7. The behaviour of power series in a boundary point of the circle of convergence (in Hungarian), Mat.Lapok, 20(1969) 347-350.

8. Matching problems. Colloquia Math. Soc. János Bolyai. 4. Combinatorial Theory and its Appl. (Balatonfüred, 1969), 695-703. (with G.O.H. Katona)

9. Poissonian random measures and linear processes with independent pieces. Bull. de l’Acad. Polonaise des Sci. Ser. Math. 18(1970) No. 8. 475-482. (with W. Woyczynski)

10. Once more on the Poisson process. Studia Sci. Math. Hung. 5(1970), 441-444.

11. The asymptotic behaviour of sums of a random number of independent random variables (in Russian). Thesis. Moscow. 1971. pp. 98.

12. Matching problems. J. of Combinatorial Theory. 10(1971), No. 1. 60-92. (with G.O.H. Katona)

13. On a problem of summation theory with random indices (in Russian). Litovski Mat. Sbornik, 11(1971), 181-187. (with B. Freyer)

14. Exercises and Problems in Probability Theory (in Hungarian) Budapest, 1971, pp. 331. (with K. Bognár, J. Mogyoródi, A. Prékopa and A. Rényi)

15. On the convergence of sums of point processes with integer marks (in Russian)., Litovski Mat. Sbornik. 11(1971), 867-874.

16. Limit theorems for stochastic processes stopped at random (in Russian). Theory of Probability and Appl. 16(1971), 557-569.

17. On the weak convergence of sequences of probability distributions (in Hungarian). Matematikai Lapok, 11(1971), 283-287.

18. On limiting classes of distributions for sums of a random number of independent, identically distributed random variables (in Russian). Theory of Probability and Appl. 17(1972), 424-439.

19. On the rate of convergence in Levy’s metric for randomly indiced sums. Colloquia Math. Soc. János Bolyai, 9(1972), 781-787.

20. Limit theorems for the distribution of the sums of a random number of random variables. Ann. of Math. Statistics, 43(1972), 1902-1913.

Publications of D. Szász 2-

21. Stability and the law of large numbers from sums of a random number of random variables. Acta Scientiarum Math. 33(1972), 269- 274.

22. On the convergence of sums of point processes with integer marks. Stochastic Point Processes, Ed. P.A.W. Lewis, Wiley, 1972, 607-615.

23. Determination of fitting discharge distribution functions (in Hungarian). Hidrologiai Közlöny, 1972, No. 1. 1-15. (with M. Domokos).

24. On rolling characteristic functions. Periodica Math. Hung. 3(1973), 13-17.

25. Limit theorems for sums of a random number of random variables. Transactions of the Sixth Prague Conf. Prague, 1973, 833-838.

26. A limit theorem for semi-Markov processes. J. of Applied Probability, 11(1974), 521-528.

27. A collision model on the two-dimensional square-lattice. Z. für Wahrscheinlichkeitstheorie, 31(1974), 75-77. (with Dao-Quang-Tuyen)

28. On a non-linear optimization problem (in Russian). Studia Sci. Math. Hung. 9(1974), 93-100.

29. On a metrization of the vague convergence. Studia Sci. Math. Hung. 9(1974), 219-222.

30. On a problem of Cox concerning controlled variability processes in R^{k} Ann. of Probability. 3(1975), 597-607. (with P. Gács)

31. Some results and problems in the limit theory of random sums. (Independent case). Colloquia Math. Soc. János Bolyai. 11(1975),

32. Shocks in a two-component paralleled system. Colloquia Math. Soc. János Bolyai. 11(1975), 347-349.

33. Particle systems with collisions. Preprint No. 26/1975 of the Math. Institute of HAS.

34. Letter to the editor: Counterexample to a theorem of D.S. Silvestrov. Theory of Probability and Appl. 20(1975), 218-219. (with P. Major)

35. Renewal theory and multicomponent reliability systems. Adv. in Applied probability. 8(1976), 239-240.

36. A problem of two lifts. Ann. of Probability. 5(1977), 550-559.

37. Uniformity in Stone’s decomposition of the renewal measure. Ann. of Probability. 5(1977), 560-564.

38. Correlation inequalities for non-purely ferromagnetic systems. J. of Statistical Physics. 19(1980), 453-459.

39. Discussion to the paper by W. Warmuth: Kritische raumlich homogene Verzweigungsprozesse mit abzahlbarer Typenmenge. Math. Nachr. 84(1978).

40. Joint diffusion on the line. J. of Statistical Physics. 23(1980), 231-240.

41. On the effect of collisions on the motion of an atom in R^{1} Ann. of Probability. 8(1980), 1968-1078. (with P. Major)

42. Random Fields. Rigorous Results in Statistical Mechanics and Quantum Field Theory. I-II. Colloquia Math. Soc. János Bolyai. Vol. 27. pp. 1111. (Co-Editors: J. Fritz and J.L. Lebowitz)

43. Dynamical theories of motion. Colloquia Math. Soc. János Bolyai, 27(1981), 1019-1031.

44. Random walk in an inhomogeneous medium with local impurities. J. of Statistical Physics. 2(1981), 527-537. (with A. Telcs)

45. Random point distributions and their applications in reliability theory and statistical physics (in Hungarian)

a) Doctor’s thesis. pp. 141. Budapest, 1981.

b) Matematikai Lapok, 30(1982), 33-57.

c) ibidem. 30(1982).

46. Ergodic theory and chaos (in Hungarian). in Chaos. Eds.: P. Szépfalusy and T. Tél, Budapest, 1982. 437-478.

47. Convergence to equilibrium of the Lorentz gas. Colloquia Math. Soc. János Bolyai, 35(1983), 757-766. (with A. Krámli)

48. Random walks with internal degrees of freedom. I. Local limit theorems. Z. für Wahrscheinlichkeitstheorie. 63*1983), 85-95. (with A. Krámli)

49. Appendix to a paper by A. Telcs entitled ”Random walks with internal states”. Colloquia Math. Soc. János Bolyai. 36(1983), 1060-1068. (with A. Telcs)

50. How to prove the CLT for the Lorentz process by using perturbation theory? Proceedings of the 3rd PSMS (1982). Akadémiai Kiadó, 1983.

51. Central limit theorem for the Lorentz process wia perturbation theory. Communications in Math. Physics. 91(1983), 519-528. (with A. Krámli)

52. Random walks with internal degrees of freedom. II. First-hitting probabilities. Z. für Wahrscheinlichkeitstheorie 68(1984), 53-64, (with A. Krámli)

53. Persistent random walks in a one-dimensional random environment. J. of Statistical Physics. 37(1984), 27-38. (with B. Tóth)

54. Level-hitting probabilities for random walks with internal states. Proceedings of the IX-th IFAC World Congress Budapest, 1984. Vol. 5, 30-35. (with A. Krámli)

55. Random walks with internal states and the Fourier law of heat conduction. Proc. of the American-Hungarian Workshop on Multivariate Analysis, ... Stanford, 1984. 28-31. (with A. Krámli and N. Simányi)

56. The problem of recurrence for Lorentz processes. Communications in Math. Physics. 98(1985), 539-552. (with A. Krámli)

57. Statistical Physics an Dynamical Systems. Rigorous Results. Progress in Physics. Vol. 10. 1985. pp. 481. Birkhauser. (Co- Editors: J. Fritz and A. Jaffe)

58. Random walks with internal degrees of freedom. III. Stationary probabilities. Probab. Th. Rel. Fields. 72(1986), 603-617. (with A. Krámli and N. Simányi)

59. Bounds for the limiting variance of the heavy particle. Communications in Math. Physics. 104(1986), 445-455. (with B. Tóth)

60. Heat conduction in caricature models of the Lorentz gas. J. of Statistical Physics, 46(1987), 303-318. (with A. Krámli and N. Simányi)

61. Towards a unified dynamical theory of the Brownian particle in an ideal gas. Communications in Mathematical Physics. 111(1987), 41- 62. (with B. Tóth)

62. A dynamical theory of the Brownian motion in the Rayleigh gas. J. of Statistical Physics. 47(1987), 681-693. (with B. Tóth)

63. A non-Wiener random walk in a 2-D Bernoulli environment. J. of Statistical Physics. 50(1988), 599-609. (with A. Krámli and P. Lukács)

64. Dispersing billiards without focal points on surfaces are ergodic. Commun. in Math. Physics. 125(1989). 439-458 (with A. Krámli and N. Simányi)

65. Ergodic properties of semi-dispersing billiards. I. Two cylindric scatteres in the 3-D torus. Nonlinearity. 2(1989), 311-326. (with A. Krámli and N. Simányi)

66. Existence of slow manifold in low order spectral models. Meteorology (to appear, with D. Dévényi, A. Krámli, T. Tél and B. Tóth)

67. Asymmetric random walks on Thue-Morse lattices. Physica D, 38(1989), 141-153 (with S. Goldstein, K. Kelly and J. Lebowitz)

68. The K-Property of Three Billiard Balls. Annals of Mathematics. 133 (1991), 37-72 (with A. Krámli and N. Simányi)

69. A ,Transversal’ Fundamental Theorem for Semi-Dispersing Billiards. Communications in Math. Physics. 129 (1990) 535-560 (with A. Krámli and N. Simányi) Erratum: ibidem 129 (1991) 207-208

70. The K-Property of Four Billiard Balls. Communications in Math. Physics. 144 (1992), 107-148 (with A. Krámli and N. Simányi)

71. Dispersing, Focusing and the Ergodicity of Billiards. in From Phase Transitions to Chaos, World Sci. Publ. ed. G. Györgyi, I. Kondor, L. Sasvári, T. Tél. 1992, 512-520

72. The K-Property of Some Planar Hyperbolic Billiards. Communications in Math. Physics. 145 (1992), 595-604

73. Ergodicity of Classical Billiard Balls. Physica A. 194 (1993) 86-92.

74. The K-Property of ‘Orthogonal’ Cylindric Billiards. Commun. Math. Phys. 160 (1994), 581-597

75. The K-Property of 4-D Billiards with Non-Orthogonal Cylindric Scatterers. J. Stat. Phys. 76 (1994) 587-604 (with N. Simányi)

76. Boltzmann’s Ergodic Hypothesis, a Conjecture for Centuries? a.) Erwin Schrödinger Institute, Vienna, Technical Report, May 1994 b.) Studia Sci. Math. Hung. 31 (1996), 299-322

77. The K-Property of Hamiltonian Systems with Restricted Hard Ball Interactions, Mathematical Research Letters, 2 (1995), 751-770 (with N. Simányi)

78. The Boltzmann-Sinai Ergodic Hypothesis for Hard Ball Systems. manuscript, 1995.

79. Hard Ball Systems are Fully Hyperbolic, Annals of Mathematics (submitted, with N. Simányi)

Scientific-Popular

1. Has the Enigma of the Margin Been Solved? (in Hungarian) Természet Világa, 124 (1993), 483-484.

2. The Role of Mathematics in Sciences. Some Thoughts about its Teaching at University. (in Hungarian) Természet Világa, 125 (1995), 514.

January, 1997